Jacob Company sold $1,000,000 of 6%, 10-year bonds at 96 on January 1, 2021. The bonds were dated January 1, 2021 and pay interest on June 30 and December 31. Jacob paid $50,000 in bond issue costs. If Jacob uses the straight-line amortization, the amount of interest expense for year 2021 would be: 

_______ 
Answer #1

Calculation of Amount of Interest Expense for Year 2021 under Straight Line Amortisation:

Interest Expense = Interest Payable on Face Value of Bond + Discounts on Bonds Payable (Straight Line Amortisation)

Semi Annual Interest Payable on Face Value of Bonds = (Face Value of Bonds * Interest Rate) * (6 Months/ 12 Months) = (6% of $10,00,000) * (6 Months / 12 Months) = $30,000

Number of Bonds Issued = $10,00,000 / $100 = 10,000

Discount on Bonds = Number of Bonds * (Face Value – Issue Price) = 10,000 * ($100 - $96) = $40,000

Discount on Bonds to be Amortised Semi Annually = (Discount on Bonds / Life of Bond) * (6 Months / 12 Months) = ($40,000 / 10 Years) * 6 Months / 12 Months = $2,000

June 30, 2021:

Interest Expense = $30,000 + $2,000 = $32,000

December 31, 2021:

Interest Expense = $30,000 + $2,000 = $32,000

Total Interest Expenses for Year 2021 = $32,000 + $32,000 = $64,000